OpenStudy (anonymous):

A physical education class has three times as many girls as boys. During a class basketball game, the girls average 18 points each, and the class as a whole averages 17 points per person. How many points does each boy score on average?

OpenStudy (anonymous):

Let n = the number of boys Then 3n = the number of girls (18 * 3n) + (x * n) = (4n * 17) 54n + x*n = 68n...........Divide both sides of this equation by n. 54 + x = 68 x = 68 - 54 = 14 The boys average 14 points.